cathecat88:

Rick and Debbie are painting a large barn. Debbie paints twice as fast as Rick. They work together for 16 hours and complete the job. Determine how fast Debbie and Rick can paint the entire barn separately

cathecat88:

Sorry, there's no diagram or measurements for this, which is why im confused.

snowflake0531:

do you know how you would set up the equation

cathecat88:

Not really.

snowflake0531:

@AZ

AZ:

Debbie paints twice as fast as Rick how would you write that as an equation? Let's call Debbie as D and Rick as R would the equation be D = 2 * R or 2 * D = R

cathecat88:

uhm.. my brain.

AZ:

just use numbers Debbie paints twice as fast as Rick so if it took Rick took 2 hours to paint Debbie is twice as fast and she could paint in 1 hour which equation makes that work if R =2 and D = 1

cathecat88:

The first one...?

AZ:

if R equals 2 and D = 1 and you said the first one D = 2 * R put the numbers in 1 = 2 * 2 does that look correct to you?

cathecat88:

uh no

cathecat88:

wait im- im so confused

AZ:

what about the second one? 2 * D = R 2 * 1 = 2

AZ:

does the second equation look correct to you?

AZ:

what's confusing you?

cathecat88:

math

cathecat88:

Yeah the second one looks like it would be correct

AZ:

okay so now we're dealing with the rates at which they complete a job so Debbie takes D hours to complete the job by herself and Rick takes R hours to complete the job by himself following along still?

cathecat88:

yes i think

AZ:

good so for these kind of questions the equation may seem weird but if you want to understand it, this website can do it much better than I can (and I don't have the time nor the patience to explain it further unless you'd like me then I can) https://www.purplemath.com/modules/workprob.htm

AZ:

but the general form of the equation is let's say we have two people working D was how many hours it takes Debbie to complete the job by herself R was how many hours it takes Rick to complete the job by herself let's say we have another variable or letter for the amount of time it takes BOTH of them to complete the job. Let's call this letter/variable T so the equation would be 1/D + 1/R = 1/T

AZ:

following along still?

cathecat88:

yeah

AZ:

Let me write it like this using the equation tool \(\dfrac{1}{\text{D}} + \dfrac{1}{\text{R}} = \dfrac{1}{\text{T}}\)

AZ:

so you can see it better now, alright good

AZ:

so remember that equation we came up with earlier? 2 * D = R

cathecat88:

mhm

AZ:

so 2D = R let's plug that into our equation with all the fractions \( \dfrac{1}{\text{D}} + \dfrac{1}{\text{R}} = \dfrac{1}{\text{T}}\) so 2D equals R so we can replace R with 2D \( \dfrac{1}{\text{D}} + \dfrac{1}{\text{2D}} = \dfrac{1}{\text{T}}\)

AZ:

Now look back at your question again It said `They work together for 16 hours and complete the job` 16 is the time it took for BOTH of them to complete the job that means 16 is our T so we have \( \dfrac{1}{\text{D}} + \dfrac{1}{\text{2D}} = \dfrac{1}{\text{16}}\)

AZ:

does all that make sense still?

cathecat88:

Yeah, I just have to read over it a couple times for me to understand it.

AZ:

once you've read it over a few times, let me know if you end up with any questions also take a look at that link I sent

AZ:

and then we can continue with finding the answer

cathecat88:

Okay, I think I'm caught up now.

AZ:

good so we had \( \dfrac{1}{\text{D}} + \dfrac{1}{\text{2D}} = \dfrac{1}{\text{16}}\) we have to solve for D do you know how to add fractions?

cathecat88:

I mean, I learned about that, but it was in like 6th grade so I pretty much forgot.

AZ:

so to add fractions, you need to have the same denominator so if you multiply the first fraction by 2 on the top and bottom, what do you get?

snowflake0531:

i suggest multiplying 1 by 2

cathecat88:

Thats 2 ;-;

snowflake0531:

yes, so what's 2 times D

cathecat88:

2D? This is so confusing

snowflake0531:

so, 2/2D

AZ:

now remember \(\dfrac{x}{z} + \dfrac{y}{z} = \dfrac{x+y}{z}\) so for our question we have \( \dfrac{2}{\text{2D}} + \dfrac{1}{\text{2D}} = \dfrac{1}{\text{16}}\) Can you add the fraction?

snowflake0531:

i recommend adding 2 to 1